OpenStudy (anonymous):

What is the exact value of 8 times q squared all over the square root of q to the seventeenth power.? Simplify if possible. Answer 8 the square root of q all over q to the sixth power. 8 the square root of q all over q to the seventh power. 8 the square root of q all over q to the fifth power. 8 the square root of q all over q to the fourth power.

OpenStudy (campbell_st):

is this the question \[\frac{8q^2}{\sqrt{q^{17}}}\]

OpenStudy (anonymous):

yup

OpenStudy (campbell_st):

ok looking at the denominator \[\sqrt{q^{17}} = \sqrt{q^{16} \times q} = \sqrt{q^{16}} \times \sqrt{q} = q^8 \] so the fraction is now \[\frac{8q^2}{q^8\sqrt{q}}\] you now need to eliminate a common factor to gte your answer

OpenStudy (campbell_st):

the fraction can be written as \[\frac{ 8 \times q \times \sqrt{q} \times \sqrt{q}}{q^8\sqrt{q}}\] eliminate common factors

OpenStudy (anonymous):

we eliminate sqrt(q)

OpenStudy (campbell_st):

yes but there is still a common factor to eliminate \[\frac{8\times q \times \sqrt{q}}{q^7\times q}\]

OpenStudy (anonymous):

oh yea q and q

OpenStudy (campbell_st):

I think you now have the answer

OpenStudy (anonymous):

is it is it with a sixth power or seventh power.

OpenStudy (campbell_st):

well you will have the 7th power in the denominator

OpenStudy (anonymous):

ok thanks